OpenStudy (anonymous):

I need help with the following please: A) Two numbers have prime factorizations of 2^2 • 3 • 5 and 2 • 3^2 • 5. Which expressions will give the greatest common factor and the least common multiple? GCF 2 • 3; LCM 2 • 3 • 5 GCF 2^2 • 3^2; LCM 2^2 • 3^2 • 5^2 GCF 2 • 3 • 5; LCM 2^2 • 3^2 • 5 GCF 2^2 • 3^2 • 5; LCM 2 • 3 thank u!!

OpenStudy (amoodarya):

great common factor = common number with minimum power gcd=gcf(3*5,3^3*5^2)=3^1 *5^1

OpenStudy (amoodarya):

2^2 • 3 • 5 and 2 • 3^2 • 5 gcd =gcf= 2^1 *3 ^1 *5^1

OpenStudy (amoodarya):

least common multiple common with maximum power * others (not common)

OpenStudy (anonymous):

thank u!!